OpenStudy (anonymous):

The lateral area of a cone is 558cm2. The radius is 31 cm. Find the slant height to the nearest tenth.

jimthompson5910 (jim_thompson5910):

Lateral Surface Area: LSA = pi*r*s 558 = pi*31*s solve for s